You worked for 3 hours on Saturday and 4 hours on Sunday. You also received a $25 tip. If you made a total of $98.50, how much do you get paid each hour?
Answer by TimothyLamb(4379)

y = m*x + 25
---
m*x = y - 25
m = (y - 25)/x
---
x = 3 + 4
x = 7
---
m = (98.50 - 25)/7
m = 10.50 dollars/hour
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---